# ResFlow: Fine-tuning Residual Optical Flow for Event-based High Temporal Resolution Motion Estimation

### v1

@@ V-C Ablation Study
We conducted comprehensive ablation studies on the DSEC-Flow dataset, with results reported in Table II. The ablation study primarily encompassed noise patterns, noise weights, velocity transformation, self-supervision, and inference frequency.
Noise Patterns. As illustrated in Fig. 3, white noise encompasses both high- and low-frequency components, while regional noise contains more low-frequency components. We propose using regional noise as the perturbation due to its similarity to the residual flow pattern. To investigate the impact of different noise patterns, we performed separate training experiments using white noise and regional noise under consistent noise intensity. As shown in Table II, regional noise significantly improves HTR residual flow prediction, supporting our hypothesis that regional noise better reflects the residual flow pattern. This observation is further supported by Fig. 6, which highlights the regional nature of residual flows. Additionally, Fig. 8 presents an example of noise correction by the residual refiner, demonstrating its ability to mitigate random perturbations and refine noisy optical flow.
Noise Weights. The weight of noise plays a critical role: excessive noise compels the residual refiner to estimate from random values, emphasizing its capability to predict the entire flow. Conversely, insufficient noise weakens the supervision, hindering the model’s ability to learn effectively. To identify the optimal configuration, we systematically adjusted the noise weight. Experimental results indicate that a noise weight of 0.3 yields the best performance.
@@ VI Conclusion
We have presented a novel residual paradigm for HTR optical flow estimation, enabling fast and accurate HTR motion estimation based on LTR linear trajectory. By decoupling LTR flow from HTR estimation, the residual paradigm mitigates the effects of event sparsity and seamlessly integrates with any LTR algorithms. Additionally, we present a regional noise strategy to facilitate the adaptation from LTR supervision to HTR inference, where regional noise effectively emulates the residual flow patterns. Residual prediction is enhanced by training the network to correct random disturbances. The noise-based strategy improves HTR performance and enables in-domain self-supervision. By predicting residual flow, our method addresses the trade-off between accuracy and frequency in HTR prediction and supports LTR supervision. Comprehensive experiments on real-world datasets validate the effectiveness and superiority of the proposed approach.

### v2

@@ V-C Ablation Study
We conducted comprehensive ablation studies on the DSEC-Flow dataset, with results reported in Table II. The ablation study primarily encompassed noise patterns, noise strength, velocity transformation, self-supervision, and inference frequency. In addition, we conducted a thorough ablation study and evaluation of the injected noise, with the results reported in Fig 9.
Noise Patterns. As illustrated in Fig. 3, white noise encompasses both high- and low-frequency components, while regional noise contains more low-frequency components. We propose using regional noise as the perturbation due to its similarity to the residual flow pattern. To investigate the impact of different noise patterns, we performed separate training experiments using white noise and regional noise under consistent noise intensity. As shown in Table II and Fig. 9, region-based noise consistently outperforms Gaussian noise in predicting HTR residual flow, supporting our hypothesis that its structured perturbation more closely aligns with the characteristics of residual flow patterns. In addition to improved accuracy, region-based noise also demonstrates greater robustness across a wide range of noise intensities. This observation is further supported by Fig. 6, which highlights the regional nature of residual flows. Additionally, Fig. 8 presents an example of noise correction by the residual refiner, demonstrating its ability to mitigate random perturbations and refine noisy optical flow.
Noise Strength. The magnitude of the injected noise plays a critical role in balancing supervision quality and learning robustness. Excessive noise forces the residual refiner to rely on near-random perturbations, effectively shifting its role toward predicting the entire flow rather than refining it. On the other hand, overly weak noise fails to introduce sufficient perturbation, leading to limited supervisory signal and suboptimal learning. To identify an effective trade-off, we systematically varied the noise strength. As shown in Table II and Fig. 9, a moderate noise weight of 0.3 achieves the best performance. Notably, compared to Gaussian noise, region-based noise demonstrates significantly greater robustness to changes in noise intensity. While the performance of Gaussian noise varies markedly with the noise level, region-based noise maintains stable and consistent improvements across a wide intensity range.
@@
Figure 9: Ablation study on noise pattern. We evaluate the impact of injected noise on model performance (FWL metric) using two types of noise: Gaussian and region-based. Each data point represents a specific noise intensity. The baseline result without noise injection is also plotted for reference.
@@
Experiments show that our residual-based framework effectively captures nonlinear motion and supports accurate HTR flow estimation. Key components such as velocity transformation, region-based noise, and self-supervision significantly enhance performance and robustness.
@@ VI Conclusion
We have presented a novel residual paradigm for HTR optical flow estimation, enabling fast and accurate HTR motion estimation based on LTR linear trajectory. By decoupling LTR flow from HTR estimation, the residual paradigm mitigates the effects of event sparsity and seamlessly integrates with any LTR algorithms. Additionally, we present a regional noise strategy to facilitate the adaptation from LTR supervision to HTR inference, where regional noise effectively emulates the residual flow patterns. Residual prediction is enhanced by training the network to correct random disturbances. The noise-based strategy improves HTR performance and enables in-domain self-supervision. In the future, we aim to reduce reliance on LTR ground truth by exploring more general self-supervised learning strategies. By predicting residual flow, our method addresses the trade-off between accuracy and frequency in HTR prediction and supports LTR supervision. Furthermore, our framework serves as a solid foundation for broader applications such as frame interpolation and high-precision motion estimation. Comprehensive experiments on real-world datasets validate the effectiveness and superiority of the proposed approach.